Grade 9Mechanics

A body travels 200 CM in the first 2 seconds and 220 CM in the next 4 seconds later same acceleration the velocity of the body at the end of 7thsecond

surendra surendra
7 Years agoGrade 9

1 Answer

Arun
7 Years ago

The average speed over the first 2 seconds is

200/2 cm/second = 100 cm/sec

The average speed over the next 4 seconds is

220/4 cm/sec or 55 cm/sec.

The body has decelerated in 3 seconds from 100 to 55 cm/sec.

That is change in speed of 100–55 = 45 cm/sec in 3 seconds. Or a deceleration is 15 cm/sec per second.

So the body is subject to a deceleration of 15 cm/sec^2.

Now we can write the speed at each second starting from zero. The body starts off at 100 + 15 = 115, and reaches 100- 15 = 85 cm/sec. at the end of the first two seconds. Over the next four seconds the speed must decrease by 4 x 15 = 60cm/sec, down to 25 cm/sec at the beginning of the seventh second. And at the end of that second will be down to 25–15 cm/sec. = 10 cm/sec.